OpenStudy (anonymous):

Prove that [sinx/1-cotx] + [cosx/1-tanx]=sinx+cosx

OpenStudy (jdoe0001):

hmmm one sec

OpenStudy (jdoe0001):

\(\bf \cfrac{sin(x)}{1-cot(x)}+\cfrac{cos(x)}{1-tan(x)}=sin(x)+cos(x) \\ \quad \\ \quad \\ \cfrac{sin(x)}{1-cot(x)}+\cfrac{cos(x)}{1-tan(x)}\implies \cfrac{sin(x)}{1-{\color{brown}{ \frac{cos(x)}{sin(x)}}}}+\cfrac{cos(x)}{1-{\color{brown}{ \frac{sin(x)}{cos(x)}}}} \\ \quad \\ \cfrac{sin(x)}{\frac{sin(x)-cos(x)}{sin(x)}}+\cfrac{cos(x)}{\frac{cos(x)-sin(x)}{cos(x)}} \\ \quad \\ \cfrac{sin(x)}{1}\cdot \cfrac{sin(x)}{sin(x)-cos(x)}+\cfrac{cos(x)}{1}\cdot \cfrac{cos(x)}{cos(x)-sin(x)} \\ \quad \\ \cfrac{sin^2(x)}{sin(x)-cos(x)}+\cfrac{cos^2(x)}{cos(x)-sin(x)} \\ \quad \\ \cfrac{sin^2(x)}{{\color{brown}{ -[cos(x)-sin(x)] }}}+\cfrac{cos^2(x)}{cos(x)-sin(x)} \\ \quad \\ \cfrac{sin^2(x)-cos^2(x)}{-[cos(x)-sin(x)]}\implies \cfrac{sin^2(x)-cos^2(x)}{sin(x)+cos(x)}\) can you see where that goes?

OpenStudy (anonymous):

Yes that is really helpful. Thank you!

OpenStudy (jdoe0001):

so.. the numerator, is just a difference of squares :) expand that, and the denominator goes kaput

OpenStudy (jdoe0001):

yw